# Processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in Inert Higgs Doublet Models and Two Higgs Doublet Models

### v2

@@ 4.2.2 Enhancement factors
The survey for all the enhancement factors at $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV are concerned in the next paragraphs. The factors $\mu_{hh}^{\textrm{THDM}}$ are large in the regions ($M_{H^{\pm}}\leq 200$ GeV) and they can reach to $1.5$. They then are decreased rapidly to the regions around $M_{H^{\pm}}\sim 300$ GeV and develop to the peak at $M_{H^{\pm}}=\sqrt{\hat{s}_{\gamma\gamma}}/2=500$ GeV. Around the peak, the enhancement factor is about $1.2$. In other ranges of charged Higgs mass, $\mu_{hh}^{\textrm{THDM}}\sim 0.9$. One also finds that $\mu_{hh}^{\textrm{THDM}}$ is inversely proportional to $t_{\beta}$ in this scenario. In $hH$ productions, the factors are increased to the peak $M_{H^{\pm}}=\sqrt{\hat{s}_{\gamma\gamma}}/2=500$ GeV and they are about $0.3$ around the peak. In all regions of $M_{H^{\pm}}$, the mentioned factors are in the ranges of [0.025,0.3][$\sim$ 0.025$\sim$ 0.3][ ∼ 0.025 , ∼ 0.3 ] and they are inversely proportional to $t_{\beta}$. In the last case, it is found that the factors $\mu_{HH}^{\textrm{THDM}}$ are same behavior as previous scenario. They are in the ranges of [0.5,4][$\sim$ 0.5$\sim$ 4][ ∼ 0.5 , ∼ 4 ] in all regions of $M_{H^{\pm}}$. However, the factors $\mu_{HH}^{\textrm{THDM}}$ depend slightly on $t_{\beta}$ in this scenario.
@@
The enhancement factors scanned over the parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ M^{2}$ in the THDM are also interested greatly in this work. Two scenarios for $c_{\beta-\alpha}>0$ and $c_{\beta-\alpha}<0$ are studied in detail in the following paragraphs. In Fig. 20 (for $c_{\beta-\alpha}>0$ scenario), Fig. 21 (for $c_{\beta-\alpha}<0$ scenario), we consider $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV (for all the above scatter plots) and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for all the below scatter plots). Moreover, we vary charged Higgs mass as $100$ GeV $\leq M_{H^{\pm}}\leq 1000$ GeV, the soft-breaking parameter as 00 GeV2 $\leq M^{2}\leq 200^{2}$ GeV2 and take $t_{\beta}=5$ for all cases.
In Fig. 20, the first scenario of $c_{\beta-\alpha}>0$ is examined. For this case, we take $c_{\beta-\alpha}=+0.2$ as an example and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, accordingly. For $hh$ production at $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV, we observe the peak of $\mu_{hh}^{\textrm{THDM}}$ at $M_{H^{\pm}}=250$ GeV which is corresponding to the threshold of cross-sections $hh$ in the THDM at the peak $\sqrt{\hat{s}_{\gamma\gamma}}=2M_{H^{\pm}}$. Around the peak, $\mu_{hh}^{\textrm{THDM}}$ are varied from $1.0$ to $1.8$. Above the peak regions, the enhancement factors tend to $1.2$ and depend slightly on $M^{2}$. For $hH$ productions, $\mu_{hH}^{\textrm{THDM}}$ are more sensitive with $M^{2}$ in below the peak regions. The factors are in the ranges of $[0.07,1.5]$ in above the peak regions. Around the peak, $\mu_{hH}^{\textrm{THDM}}$ can reach to $0.3$. We also find the same behavior for $\mu_{HH}^{\textrm{THDM}}$. The factors for $HH$ productions are large in the low regions of $M_{H^{\pm}}$ and around the peak $M_{H^{\pm}}=250$ GeV. They are in the ranges of $[0.7,2.4]$ in the above the peak regions. Generally, we observe that $\mu_{\phi_{i}\phi_{j}}^{\textrm{THDM}}$ are proportional to $1/M$ at this CoM energy.
At $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV, we also find that $\mu_{hh}^{\textrm{THDM}}$ develop to the peak at $M_{H^{\pm}}=500$ GeV where the factors can reach to $2.2$ and are decreased rapidly beyond the peak. The factors depend slightly on $M^{2}$ and tend to $1$ beyond the peak regions. For $hH$ productions, $\mu_{hH}^{\textrm{THDM}}$ are sensitive with $M^{2}$ in the peak regions. They tend to $0.05$ and they are slightly dependent of $M^{2}$ in the above the peak regions. For $HH$ productions, the factors become large in the below the peak regions and they are inversely proportional to $M^{-1}$. Around the peak, the factors are enhanced by large values of $M^{2}$. Above the peak regions, $\mu_{HH}^{\textrm{THDM}}$ are varied around $1.0$.
@@
We next comment on physical results at $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV. The factors $\mu_{hh}^{\textrm{THDM}}$ are enhanced by the small values of $M$ in low regions of charged Higgs mass and they can reach $1.2$. They tend to $2.5$ around the peak. The factors are then varied around $\sim 0.9$. In general, the factors depend on $M^{-1}$. In the productions $hH$, $\mu_{hH}^{\textrm{THDM}}$ are more sensitive to $M^{-1}$ around the peak regions. They then tend to $0.2$ in the high mass regions of singly charged Higgs. For $HH$ productions, the factors $\mu_{HH}^{\textrm{THDM}}$ strongly depend on $M^{-1}$. At the peak, the factors are enhanced by the large value of $M$. Above the peak regions, $\mu_{HH}^{\textrm{THDM}}$ tend to $\sim 1$.
@@ 5 Conclusions
In this paper, we have presented the results for one-loop induced processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,\leavevmode\nobreak\ H$ at high energy photon-photon collision in the IHDM and the THDM. In the phenomenological results, we have shown the cross-sections at several center-of-mass energies. The results show that cross-sections for the computed processes in the models under investigations are enhanced at around the threshold of charged Higgs pair ($\sim 2M_{H^{\pm}}$). Furthermore, the enhancement factors for the processes are examined in parameter space of the models under consideration. In the IHDM, the factors are studied in the parameter space of $(M_{H^{\pm}},\leavevmode\nobreak\ \mu^{2}_{2})$ and $(M_{H^{\pm}},\leavevmode\nobreak\ \lambda_{2})$. In the the THDM, the factors are analysed in the planes of $(M_{H^{\pm}},\leavevmode\nobreak\ t_{\beta})$ and $(M_{H^{\pm}},\leavevmode\nobreak\ M^{2})$. Two scenarios of $c_{\beta-\alpha}>0$ and $c_{\beta-\alpha}<0$ have studied in further detail. The factors give a different behavior from considering these scenarios. As a result, discriminations for the above-mentioned scenarios can be performed at future colliders.

### v3

@@ 3.2.2 Enhancement factors
The survey for all the enhancement factors at $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV is presented in the next paragraphs. The factors $\mu_{hh}^{\textrm{THDM}}$ are large in the regions ($M_{H^{\pm}}\leq 200$ GeV) and they can reach up to $1.5$. They then decrease rapidly to the regions around $M_{H^{\pm}}\sim 300$ GeV and develop a peak at $M_{H^{\pm}}=\sqrt{\hat{s}_{\gamma\gamma}}/2=500$ GeV. Around the peak, the enhancement factor is about $1.2$. In other ranges of charged Higgs mass, $\mu_{hh}^{\textrm{THDM}}\sim 0.9$. One also observes that $\mu_{hh}^{\textrm{THDM}}$ is inversely proportional to $t_{\beta}$ in this scenario. In $hH$ productions, the factors increase to the peak at $M_{H^{\pm}}=\sqrt{\hat{s}_{\gamma\gamma}}/2=500$ GeV and they are about $0.3$ around the peak. In all regions of $M_{H^{\pm}}$, the mentioned factors are in the ranges of [0.025,0.3][$\sim$ 0.025$\sim$ 0.3][ ∼ 0.025 , ∼ 0.3 ] and they are inversely proportional to $t_{\beta}$. In the last case, it is found that the factors $\mu_{HH}^{\textrm{THDM}}$ show the same behavior as the previous scenario. They are in the ranges of [0.5,4][$\sim$ 0.5$\sim$ 4][ ∼ 0.5 , ∼ 4 ] in all regions of $M_{H^{\pm}}$. However, the factors $\mu_{HH}^{\textrm{THDM}}$ depend only slightly on $t_{\beta}$ in this scenario.
@@
[…]
In Fig. 7, the first scenario of $c_{\beta-\alpha}>0$ is examined. For this case, we take $c_{\beta-\alpha}=+0.2$ as an example and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, accordingly. For $hh$ production at $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV, we observe the peak of $\mu_{hh}^{\textrm{THDM}}$ at $M_{H^{\pm}}=250$ GeV, which is corresponding to the threshold of cross-sections for $hh$ in the THDM at the peak $\sqrt{\hat{s}_{\gamma\gamma}}=2M_{H^{\pm}}$. Around the peak, $\mu_{hh}^{\textrm{THDM}}$ varies from $1.0$ to $1.8$. Above the peak regions, the enhancement factors tend to $1.2$ and depend slightly on $M^{2}$. For $hH$ production, $\mu_{hH}^{\textrm{THDM}}$ is more sensitive to $M^{2}$ below the peak regions. The factors are in the range of $[0.07,1.5]$ above the peak regions. Around the peak, $\mu_{hH}^{\textrm{THDM}}$ can reach up to $0.3$. We also find the same behavior for $\mu_{HH}^{\textrm{THDM}}$. The factors for $HH$ production are large in the low regions of $M_{H^{\pm}}$ and around the peak $M_{H^{\pm}}=250$ GeV. They are in the range of $[0.7,2.4]$ above the peak regions. Generally, we observe that $\mu_{\phi_{i}\phi_{j}}^{\textrm{THDM}}$ is proportional to $1/M$ at this CoM energy.
At $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV, we also find that $\mu_{hh}^{\textrm{THDM}}$ develops to the peak at $M_{H^{\pm}}=500$ GeV, where the factors can reach up to $2.2$ and decrease rapidly beyond the peak. The factors depend slightly on $M^{2}$ and tend to $1$ beyond the peak regions. For $hH$ production, $\mu_{hH}^{\textrm{THDM}}$ is sensitive to $M^{2}$ in the peak regions. They tend to $0.05$ and are slightly dependent on $M^{2}$ in the regions above the peak. For $HH$ production, the factors become large in the regions below the peak and they are inversely proportional to $M^{-1}$. Around the peak, the factors are enhanced by large values of $M^{2}$. Above the peak regions, $\mu_{HH}^{\textrm{THDM}}$ varies around $1.0$.
@@
We next comment on the physical results at $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV. The factors $\mu_{hh}^{\textrm{THDM}}$ are enhanced by the small values of $M$ in the low regions of charged Higgs mass and they can reach up to $1.2$. They tend to $2.5$ around the peak. The factors are then varied around $\sim 0.9$. In general, the factors depend on $M^{-1}$. In the $hH$ production, $\mu_{hH}^{\textrm{THDM}}$ is more sensitive to $M^{-1}$ around the peak regions. They then tend to $0.2$ in the high mass regions of singly charged Higgs. For $HH$ production, the factors $\mu_{HH}^{\textrm{THDM}}$ strongly depend on $M^{-1}$. At the peak, the factors are enhanced by the large value of $M$. Above the peak regions, $\mu_{HH}^{\textrm{THDM}}$ tend to $\sim 1$.
@@ 4 Conclusions
In this paper, we have presented the phenomenological results for one-loop induced processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,\leavevmode\nobreak\ H$ at high-energy photon-photon collisions in the IHDM and the THDM. In the phenomenological results, we have shown the cross sections at several center-of-mass energies. The results show that cross sections for the computed processes in the models under investigation are enhanced around the threshold of charged Higgs pair production ($\sim 2M_{H^{\pm}}$). Furthermore, the enhancement factors for the processes are examined in the parameter space of the models under consideration. In the IHDM, the factors are studied in the parameter space of $(M_{H^{\pm}},\leavevmode\nobreak\ \mu_{2}^{2})$ and $(M_{H^{\pm}},\leavevmode\nobreak\ \lambda_{2})$. In the THDM, the factors are analyzed in the planes of $(M_{H^{\pm}},\leavevmode\nobreak\ t_{\beta})$ and $(M_{H^{\pm}},\leavevmode\nobreak\ M^{2})$. Two scenarios of $c_{\beta-\alpha}>0$ and $c_{\beta-\alpha}<0$ have been studied in further detail. The factors give a different behavior when considering these scenarios. As a result, discriminations for the above-mentioned scenarios can be performed at future colliders. The dependence of the cross section for the process $\gamma\gamma\rightarrow hH$ on $m_{12}^{2}$ provides a potential probe of the soft $Z_{2}$-breaking scale in the Two Higgs Doublet Model.